# Anchored Supervised Fine-Tuning

### v2

@@ 6.4 Cross-Domain Validation on Code Generation
Methods HumanEval (%) HumanEval+ (%) MBPP (%) MBPP+ (%) Avg (%)
LLamMA2-7B 17.1 14.6 28.3 21.2 20.3
SFT 23.2 19.5 34.1 28.6 26.4
iw-SFT 23.8 21.3 31.0 26.5 25.7
DFT 15.9 12.8 28.3 22.3 19.8
[…]
Table 3: Performance on code generation with LLaMA-2-7B.
@@ 6.5 Computational Efficiency and Memory Analysis
To address these practical limitations, we propose ASFT-LoRA, leveraging the mathematical properties of low-rank adaptation to enable memory-efficient implementation. The key insight exploits LoRA’s parameter decomposition $\Delta W=BA$ where the fine-tuned model becomes $\pi_{\theta}(y|x)=\pi_{\mathrm{base}}(y|x;W_{\mathrm{base}}+BA)$. This parameterization enables computing $D_{\mathrm{KL}}(\pi_{\theta}(\cdot|s)|\pi_{\mathrm{base}}(\cdot|s))$ using a single model instance by dynamically switching between $W_{\mathrm{base}}$ and $W_{\mathrm{base}}+BA$ during forward passes, eliminating the need for separate model copies while preserving theoretical anchoring guarantees. As shown in Table 4, ASFT-LoRA (rank = 8) reduces memory consumption to 40.70GB (comparable to standard SFT) and training time to 0.562 hours (7.3% overhead over SFT), while maintaining meaningful improvements over SFT baselines (34.32 vs 33.04 overall accuracy). Although performance degrades compared to full-parameter ASFT (42.03 overall accuracy), ASFT-LoRA provides a practical compromise for resource-constrained environments while demonstrating the versatility of our anchoring framework across different parameter-efficiency regimes.
@@ 7 Conclusion
We provide comprehensive implementation details in Section 5, including specific hyperparameters, datasets, and evaluation protocols. All experiments use publicly available datasets and standard benchmarks. We will release our implementation code to ensure reproducibility of our results.
@@ D.1 Derivation of SFT as RL Lower Bound
We provide a complete proof of Proposition 1, following the theoretical framework established in (Qin & Springenberg, 2025), showing that SFT optimizes a lower bound on the RL objective in sparse reward settings.
@@
D.2 From Importance Sampling to Importance-Weighted Lower Bounds
@@
[Importance-Weighted Lower Bound] Starting from the importance-sampled RL objective:
@@
The first and third terms are independent of $\theta$, so we can drop them from the optimization objective:
@@
D.3 From Importance Sampling to Reward-Weighted Lower Bounds
We now derive the generalized reward-weighted framework that enables tighter bounds through auxiliary distributions.
Appendix E Proof of Tightness
@@
Appendix F Analysis
F.1 Model Scale Results
Model Method MedQA (%) MMLU (%) MedMCQA (%) Average (%)
@@ F.1 Model Scale Results
SFT 33.31 (+3.46) 33.52 (+3.00) 33.28 (-0.48) 33.37 (+1.99)
SFT w KL 29.22 (-0.63) 30.63 (+0.11) 33.01 (-0.75) 30.95 (-0.43)
iw-SFT 35.35 (+5.50) 38.92 (+8.40) 34.74 (+0.98) 36.34 (+4.96)
DFT 29.69 (-0.16) 26.69 (-3.83) 31.20 (-2.56) 29.19 (-2.19)
GRPO 30.48 (+0.63) 32.46 (+1.94) 34.64 (+0.88) 32.53 (+1.15)
DAPO 39.75 (+9.90) 48.63 (+18.11) 38.37 (+4.61) 42.25 (+10.87)
[…]
@@
DFT 36.84 (-10.53) 46.77 (-18.55) 42.39 (-4.82) 42.00 (-11.30)
SFT 41.24 (-6.13) 47.02 (-18.30) 36.43 (-10.78) 41.56 (-11.74)
[…]
@@
SFT 59.07 (-3.46) 71.19 (-3.65) 49.80 (-8.65) 60.02 (-5.25)
SFT w KL 61.12 (-1.41) 69.77 (-5.07) 51.52 (-6.93) 60.80 (-4.47)
DFT 36.45 (-26.08) 54.65 (-20.19) 45.90 (-12.55) 45.67 (-19.60)
GRPO 63.00 (+0.47) 76.16 (+1.32) 58.93 (+0.48) 66.03 (+0.76)
DAPO 63.94 (+1.41) 73.82 (-1.02) 58.88 (+0.43) 65.55 (+0.28)
[…]
@@
SFT 67.09 (+5.19) 76.12 (+13.62) 61.20 (+16.11) 68.14 (+11.64)
DFT 68.42 (+6.52) 57.03 (-5.47) 42.51 (-2.58) 55.99 (-0.51)
[…]
@@
SFT 28.04 (-47.53) 34.57 (-43.71) 38.92 (-25.99) 33.84 (-39.08)
DFT 71.01 (-4.56) 79.04 (+0.76) 64.69 (-0.22) 71.58 (-1.34)
[…]
Table 5: Performance of various fine-tuning methods on medical domain datasets for both LLaMA-2 and Qwen2.5 series. Numbers are expressed as percentages. Bold numbers indicate the best performance in each column, while rows with blue background highlight our ASFT approach. (olive) and (red) show the difference compared to each series’ Base model.
F.2 Token Probability Distribution Analysis
Figure 4: Token probability distribution across different models on the numina_cot training dataset. The x-axis shows the probability intervals of generated tokens, while the y-axis reports the percentage of tokens falling into each interval on a logarithmic scale.
Appendix G Abaltion Study
G.1 Training Hyper-Parameters
@@
Appendix H Dataset Example
H.1 Medical Knowledge Datasets
@@ A. Impulse through myelinated fibers is slower than non-myelinated fibers
H.2 Math Reasoning Datasets

### v3

@@ 6.4 Cross-Domain Validation on Code Generation
Methods HumanEval HumanEval+ MBPP MBPP+ Avg.
LLaMA-2-7B 17.1 14.6 28.3 21.2 20.3
SFT 23.2 $\uparrow$6.1 19.5 $\uparrow$4.9 34.1 $\uparrow$5.8 28.6 $\uparrow$7.4 26.4 $\uparrow$6.1
iw-SFT 23.8 $\uparrow$6.7 21.3 $\uparrow$6.7 31.0 $\uparrow$2.7 26.5 $\uparrow$5.3 25.7 $\uparrow$5.4
DFT 15.9 $\downarrow$1.2 12.8 $\downarrow$1.8 28.3 $\pm$0.0 22.3 $\uparrow$1.1 19.8 $\downarrow$0.5
ASFT 27.2 $\uparrow$10.1 21.4 $\uparrow$6.8 32.5 $\uparrow$4.2 26.7 $\uparrow$5.5 27.0 $\uparrow$6.7
Table 3: Performance (%) on code generation with LLaMA-2-7B. Bold numbers indicate the best performance in each column, and rows with blue background highlight our ASFT approach.
@@ 6.5 Computational Efficiency and Memory Analysis
To address these practical limitations, we propose ASFT-LoRA, leveraging the mathematical properties of low-rank adaptation to enable memory-efficient implementation. The key insight exploits LoRA’s parameter decomposition $\Delta W=BA$ where the fine-tuned model becomes $\pi_{\theta}(y|x)=\pi_{\mathrm{base}}(y|x;W_{\mathrm{base}}+BA)$. This parameterization enables computing $D_{\mathrm{KL}}(\pi_{\mathrm{base}}(\cdot|s)\|\pi_{\theta}(\cdot|s))$ using a single model instance by dynamically switching between $W_{\mathrm{base}}$ and $W_{\mathrm{base}}+BA$ during forward passes, eliminating the need for separate model copies while preserving theoretical anchoring guarantees. As shown in Table 4, ASFT-LoRA (rank=8, lr=5e-4) reduces memory consumption to 40.70GB (comparable to standard SFT) and training time to 0.594 hours (13.4% overhead over SFT), while achieving substantial improvements over SFT baselines (39.45 vs 33.04 overall accuracy). Notably, ASFT-LoRA with appropriate learning rate selection achieves 93.9% of full-parameter ASFT’s performance (39.45 vs 42.03 overall accuracy) while requiring less than half the memory, demonstrating the versatility of our anchoring framework across different parameter-efficiency regimes. We observe that LoRA-based training benefits from higher learning rates (5e-4) compared to full fine-tuning (2e-5), consistent with recent findings on low-rank adaptation (Hu et al., 2021).
6.6 Alternative Anchoring Mechanisms.
Our theoretical framework reveals that the key to stable optimization lies in providing distributional anchoring to prevent drift, and in principle, any regularization term that constrains the policy distribution can serve this purpose. While KL divergence provides effective anchoring, we investigate whether other terms—particularly the standard SFT loss itself—can offer a more lightweight alternative. The SFT loss $\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}[\log\pi_{\theta}(y^{*}\mid x)]$ naturally encourages the model to stay close to demonstrated behaviors, suggesting it may provide implicit anchoring effects as observed in Figure 1.
We therefore explore a hybrid objective that combines DFT’s adaptive reweighting with SFT-based anchoring:
$\mathcal{L}_{\mathrm{ASFT-SFT}}(\theta)=\mathcal{L}_{\mathrm{DFT}}(\theta)+\alpha\mathcal{L}_{\mathrm{SFT}}(\theta)$ (7)
where $\alpha$ controls the anchoring strength. Compared to KL-based anchoring, this approach is computationally more efficient as it does not require maintaining a separate reference model during training. As shown in Figure 4, we conduct a grid search over $\alpha\in\{0.01,0.05,0.1,0.5,1.0,5.0,10.0\}$ on both math and medical tasks.
Interestingly, we observe domain-specific patterns: on mathematical reasoning tasks, ASFT-SFT achieves competitive performance at moderate $\alpha$ values (0.05-0.5), but degrades significantly at larger coefficients (dropping from 0.275 to 0.168 as $\alpha$ increases from 0.01 to 10.0). This suggests that for reasoning tasks where the base model already possesses strong prior knowledge, moderate SFT anchoring can effectively stabilize training by activating these capabilities without over-constraining exploration. In contrast, on medical knowledge tasks, ASFT-SFT shows more volatile behavior across different $\alpha$ values, with performance fluctuating between 0.340 and 0.370. This instability likely reflects the challenge of balancing knowledge injection (requiring departure from the base distribution) with stability preservation in knowledge-intensive domains. Throughout both settings, ASFT with KL anchoring (green lines) maintains consistently superior and stable performance across all coefficient values, demonstrating that while SFT loss provides useful anchoring, KL divergence offers more robust distributional control, particularly for tasks requiring significant distribution shift from the base model.
Figure 4: Comparison of KL-based anchoring (ASFTKL) versus SFT loss-based anchoring (ASFTSFT) across different coefficient values ($\alpha$ or $\lambda$) on math reasoning and medical knowledge tasks. ASFTKL maintains stable performance across all values, while ASFTSFT shows domain-specific sensitivity: degradation at large coefficients for math tasks and volatility for medical tasks.
@@ 7 Conclusion
Acknowledgements
This project was supported by National Natural Science Foundation of China (No. 62306132), Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515011564), Natural Science Foundation of Shanghai (No. 25ZR1402136). We thank the anonymous reviewers for their insightful feedback on this work.
@@ Acknowledgements
We provide comprehensive implementation details in Section 5, including specific hyperparameters, datasets, and evaluation protocols. All experiments use publicly available datasets and standard benchmarks. We have released our implementation code to ensure reproducibility of our results. The code is available at https://github.com/zhuchichi56/ASFT.
Limitations
Our theoretical analysis shows that DFT’s effectiveness is fundamentally dependent on the prior knowledge embedded in the base model, which accounts for its strong performance in reasoning tasks with well-established priors, but weaker results in knowledge-intensive domains requiring more substantial distributional shifts. Although our KL-based anchoring mitigates DFT’s instability, the method’s reliance on model priors persists. Our ablation study on alternative anchoring mechanisms (Section 6.6) suggests that domain-adaptive anchoring strategies could enable more robust generalization across domains; however, developing a systematic framework for such adaptation remains an open challenge.
Additionally, ASFT incurs higher computational costs than standard SFT due to the KL divergence computation, requiring doubled GPU memory (from 38.96GB to 88.02GB for LLaMA-2-7B) and  23.7% training time overhead. Our ASFT-LoRA variant (Section 6.5) provides a practical solution by reducing memory to 40.70GB with only 7.3% overhead, though with some performance trade-off.
Future Work
A promising direction for future research is extending ASFT to continual learning settings, where models must adapt to new tasks while preserving previously acquired knowledge. The anchoring mechanism in ASFT naturally aligns with continual learning goals: the KL regularization that prevents distributional drift during single-task fine-tuning could be adapted to prevent catastrophic forgetting when learning sequential tasks. By maintaining anchors to both the base model and checkpoints from previous tasks, ASFT’s framework could potentially enable stable, progressive capability acquisition across multiple domains.
@@ D.1 Derivation of SFT as RL Lower Bound
We provide a complete proof of Proposition 1, following the theoretical framework established in (Qin and Springenberg, 2025), showing that SFT optimizes a lower bound on the RL objective in sparse reward settings.
@@
D.2 Frozen Auxiliary Distribution and Optimization Validity
We formalize how DFT’s stop-gradient mechanism ensures valid lower bound optimization by treating the auxiliary distribution as constant during each update step.
Lemma 1 (Frozen-q Surrogate).
At iteration $k$, define the auxiliary distribution
$q_{k}(\tau)\propto\pi_{\mathrm{ref}}(\tau\mid D^{+})\,p_{\theta_{k}}(\tau)$ (17)
where $\theta_{k}$ is the current parameter. Then optimizing
$\max_{\theta}\mathbb{E}_{\tau\in D^{+}}\left[\frac{q_{k}(\tau)}{\pi_{\mathrm{ref}}(\tau)}\log\pi_{\theta}(\tau)\right]$ (18)
with $q_{k}$ treated as constant (independent of $\theta$) yields a valid lower bound on $J(\theta)$ that differs only by a $\theta$-independent constant.
Proof.
Following the derivation in Appendix D.3, we have:
$J(\theta)\geq\mathbb{E}_{\tau\sim\pi_{\mathrm{ref}}}\left[\frac{q_{k}(\tau)}{\pi_{\mathrm{ref}}(\tau)}R(\tau)\right]+\mathbb{E}_{\tau\sim\pi_{\mathrm{ref}}}\left[\frac{q_{k}(\tau)}{\pi_{\mathrm{ref}}(\tau)}R(\tau)\log\pi_{\theta}(\tau)\right]\\ -\mathbb{E}_{\tau\sim\pi_{\mathrm{ref}}}\left[\frac{q_{k}(\tau)}{\pi_{\mathrm{ref}}(\tau)}R(\tau)\log q_{k}(\tau)\right]$ (19)
When $q_{k}$ is constructed from frozen parameter $\theta_{k}$ (via stop-gradient), the first and third terms become constants independent of the optimization variable $\theta$. Therefore, maximizing the second term is equivalent to maximizing a valid lower bound on $J(\theta)$.
The stop-gradient operator $\operatorname{sg}[\cdot]$ in DFT implements this frozen-$q$ mechanism: at each iteration, $q_{k}$ depends on $\theta_{k}$ from the previous step, but is treated as fixed when computing gradients with respect to the current $\theta$. ∎
This formalization clarifies that the ”dropped constant terms” in our derivations are legitimate because they are genuinely $\theta$-independent within each optimization iteration, making DFT a valid iterative lower bound maximization procedure.
D.3 From Importance Sampling to Importance-Weighted Lower Bounds
@@
Starting from the importance-sampled RL objective:
@@
When treating $q$ as fixed (independent of $\theta$, see Lemma 1), the first and third terms are constants, so we can drop them from the optimization objective:
@@
D.4 Proof of Tightness
@@
Appendix E Experiments
E.1 Model Scale Results
Model Methods MedQA MMLU MedMCQA Avg.
@@ E.1 Model Scale Results
SFT 33.31 $\uparrow$3.46 33.52 $\uparrow$3.00 33.28 $\downarrow$0.48 33.37 $\uparrow$1.99
SFT w KL 29.22 $\downarrow$0.63 30.63 $\uparrow$0.11 33.01 $\downarrow$0.75 30.95 $\downarrow$0.43
iw-SFT 35.35 $\uparrow$5.50 38.92 $\uparrow$8.40 34.74 $\uparrow$0.98 36.34 $\uparrow$4.96
DFT 29.69 $\downarrow$0.16 26.69 $\downarrow$3.83 31.20 $\downarrow$2.56 29.19 $\downarrow$2.19
GRPO 30.48 $\uparrow$0.63 32.46 $\uparrow$1.94 34.64 $\uparrow$0.88 32.53 $\uparrow$1.15
DAPO 39.75 $\uparrow$9.90 48.63 $\uparrow$18.11 38.37 $\uparrow$4.61 42.25 $\uparrow$10.87
ASFT 39.28 $\uparrow$9.43 46.37 $\uparrow$15.85 40.45 $\uparrow$6.69 42.03 $\uparrow$10.65
@@
DFT 36.84 $\downarrow$10.53 46.77 $\downarrow$18.55 42.39 $\downarrow$4.82 42.00 $\downarrow$11.30
SFT 41.24 $\downarrow$6.13 47.02 $\downarrow$18.30 36.43 $\downarrow$10.78 41.56 $\downarrow$11.74
ASFT 49.57 $\uparrow$2.20 65.86 $\uparrow$0.54 50.61 $\uparrow$3.40 55.35 $\uparrow$2.05
@@
SFT 59.07 $\downarrow$3.46 71.19 $\downarrow$3.65 49.80 $\downarrow$8.65 60.02 $\downarrow$5.25
SFT w KL 61.12 $\downarrow$1.41 69.77 $\downarrow$5.07 51.52 $\downarrow$6.93 60.80 $\downarrow$4.47
DFT 36.45 $\downarrow$26.08 54.65 $\downarrow$20.19 45.90 $\downarrow$12.55 45.67 $\downarrow$19.60
GRPO 63.00 $\uparrow$0.47 76.16 $\uparrow$1.32 58.93 $\uparrow$0.48 66.03 $\uparrow$0.76
DAPO 63.94 $\uparrow$1.41 73.82 $\downarrow$1.02 58.88 $\uparrow$0.43 65.55 $\uparrow$0.28
ASFT 61.98 $\downarrow$0.55 76.60 $\uparrow$1.76 59.00 $\uparrow$0.55 65.86 $\uparrow$0.59
@@
SFT 67.09 $\uparrow$5.19 76.12 $\uparrow$13.62 61.20 $\uparrow$16.11 68.14 $\uparrow$11.64
DFT 68.42 $\uparrow$6.52 57.03 $\downarrow$5.47 42.51 $\downarrow$2.58 55.99 $\downarrow$0.51
ASFT 71.80 $\uparrow$9.90 79.23 $\uparrow$16.73 58.76 $\uparrow$13.67 69.93 $\uparrow$13.43
@@
SFT 28.04 $\downarrow$47.53 34.57 $\downarrow$43.71 38.92 $\downarrow$25.99 33.84 $\downarrow$39.08
DFT 71.01 $\downarrow$4.56 79.04 $\uparrow$0.76 64.69 $\downarrow$0.22 71.58 $\downarrow$1.34
ASFT 77.69 $\uparrow$2.12 87.08 $\uparrow$8.80 71.00 $\uparrow$6.09 78.59 $\uparrow$5.67
Table 5: Performance of various fine-tuning methods on medical domain datasets for both LLaMA-2 and Qwen2.5 series. Bold numbers indicate the best performance in each group, and rows with blue background highlight our ASFT approach. Arrows with $\uparrow$ and $\downarrow$ indicate performance improvements and degradations relative to the base model. All values are reported in percentages.
E.2 Training Hyper-Parameters
@@
Appendix F Dataset Example
F.1 Medical Knowledge Datasets
@@ A. Impulse through myelinated fibers is slower than non-myelinated fibers
F.2 Math Reasoning Datasets